OpenStudy (anonymous):

If cos t = 3/5 and 3π/2 < t < 2π, compute sin t and cot t.

OpenStudy (anonymous):

Alright. First things first: I'm more comfortable with degrees as opposed to radians. Are you the same way?

OpenStudy (anonymous):

I am not sure what they are asking me to find.

OpenStudy (anonymous):

They want you to solve the sin of t and the cotangent of t.

OpenStudy (anonymous):

The reason they gave us that scary 3pi/2 junk is to tell us in which quadrant our triangle will be.

OpenStudy (anonymous):

Do you have about 5 minutes free to dedicate to this question?

OpenStudy (anonymous):

i see so I would just to create a triangle and find the sin and cotangent

OpenStudy (anonymous):

Yup. BUT.. we have to know which quadrant to make it. It's going to take about five minutes to explain it, so lemme know if you have that kind of time available.

OpenStudy (anonymous):

I'll take that as a no, then?

OpenStudy (anonymous):

I was following along uber...is the sin t = -4/5 ?

OpenStudy (anonymous):

Oh. I was hoping to create a scribbla page. Just gie me like 5 minutes to help someone else with her question, and I'll get back with you.

OpenStudy (anonymous):

Actually, I should be able to multitask. Could you meet me in here: http://www.scribblar.com/u306z82j

OpenStudy (anonymous):

You are told that \(\cos t =\frac{3}{5}\), and that it is for values of t, between \(3 \pi/2<t<2\pi\). This means that t is an angle between these values, in the fourth quadrant (see attached picture). Firstly, to find \(\sin t\), use the relationship (which is always true) that \[\sin^2 t + \cos^2 t = 1\] In other words \[\sin^2 t + \left(\frac{3}{5}\right)^2=1\] so we have that \[\sin^2 t = 1 - \left(\frac{9}{25}\right)=\frac{16}{25}\] so we have that \(\sin t = + \frac{4}{5}\) OR \(\sin t =- \frac{4}{5}\). This is where the values of t come in. As only those of the fourth quadrant are allowed, the sine function is negative, and so we take that \[\sin t = - \frac{4}{5}\] Secondly, to get \(\cot t\), we can use the relationship that \[\tan \theta = \frac{\sin \theta}{\cos \theta}\] and that \(\cot \theta = \frac{1}{\tan \theta}\) so you have, \[\tan t = \frac{\sin t }{\cos t }= \frac{-4/5}{3/5}= \frac{-4}{3}\] and then \[\cot t = \frac{1}{tan t}= -\frac{3}{4}\] Hope that helps

OpenStudy (anonymous):

Opps here is the figure